Convert the given amounts to the respective blank cells. Fill in ALL the blanks in the table below. Round each to the nearest CENT.

Annual

Monthly

Bi-Weekly

Weekly

Hourly

41,579

4,073

780

1,472
Answer #1

Answer

A B=A/12 C=A/26 D=A/52 E=A/(365 x 24)
Annual Monthly Bi-Weekly Weekly Hourly
$    41,579.00 $       3,464.92 $       1,599.19 $          799.60 $                     4.75
$    48,876.00 $       4,073.00 $       1,879.85 $          939.92 $                     5.58
$    20,280.00 $       1,690.00 $          780.00 $          390.00 $                     2.32
$    76,544.00 $       6,378.67 $       2,944.00 $       1,472.00 $                     8.74